Three friends want to go camping and they pack enough water to last for 6 days. Just as they are about to leave they are joined by 3 more people. How long will their water last now?


1
Expert's answer
2021-10-04T19:28:14-0400

"\frac{W}{3}=6 \to W=6*3=18."

"\frac{W}{3+3}=\frac{18}{6}=3."

The water lasts for 3 days.